OpenStudy (anonymous):

verify this identity tan (x+pi/2)= -cotx

OpenStudy (anonymous):

tan(x+pi/2) = sin(x+pi/2)/cos(x+pi/2) = cos(x)/-sin(x) = -cot(x)

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

sin(x+pi/2) = cos(x) and cos(x+pi/2) = -sin(x)

OpenStudy (anonymous):

@$w3G_Godd how did you get cos(x)/-sin(x)?

OpenStudy (anonymous):

tan(anything) = sin(anything)/cos(anything) so tan(2) = sin(2)/cos(2) tan(2x) = sin(2x)/cos(2x) tan(puppy dog) = sin(puppy dog)/cos(puppy dog) so tan(x+pi/2) = sin(x+pi/2)/cos(x+pi/2)

OpenStudy (anonymous):

better?

OpenStudy (anonymous):

ya i got that, but why does it then flip to cos(x)/-sin(x)?

OpenStudy (anonymous):

sin(x+pi/2) = sin(x)cos(pi/2) + cos(x)sin(pi/2) = 0+cos(x)=cos(x) cos(x+pi/2)= cos(x)cos(pi/2) - sin(x)sin(pi/2)= 0-sin(x)=-sin(x)

OpenStudy (anonymous):

sorry i just now saw this